OpenStudy (anonymous):

Find the 18th term of the sequence 5, 8, 11, 14, 17... a. An 18th term does not exist b. 53 c. 56

Parth (parthkohli):

If you're gonna stay on this site for a while, then I suggest you to make a new account with a different username C:

Parth (parthkohli):

By the way, the \(n\) term of an arithmetic sequence is \(a_1 \times (n - 1)d\)

OpenStudy (anonymous):

What does d mean?

Parth (parthkohli):

The common difference is \(d\).

Parth (parthkohli):

You just subtract the second term from the first term, and you get \(d\). Here, \(d = 8 - 5\)

OpenStudy (anonymous):

What's a1?

Parth (parthkohli):

The first term. Here, it is \(3\).

OpenStudy (anonymous):

The first term is the same as the common difference? How did you get 3?

OpenStudy (anonymous):

COME BACK

OpenStudy (anonymous):

is it 53?

OpenStudy (anonymous):

I need serious help

OpenStudy (anonymous):

general formula is \[a_{n}=3n+2\] when n=18 \[a_{18}=3(18)+2=56\]